Complete the table of ordered pairs for the given linear equation x=-y+3
Answer by ewatrrr(24785)

 
Hi,
Complete the table of ordered pairs for the given linear equation
x= -y+3 0r y = -x + 3
x y = -x +3
-2 5
-1 4
0 3
1 2
2 1